Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 126     Entry time: 14 Jun 2024, 01:53
Author: Jörg Hallmann 
Group: MID 
Subject: Shift summary 

continue with Tempus tests
continue with PAM commissioning (first results of the working device at MID)

comments:
changed to 10 keV in order to use only one focussing lens in CRL2 and generating a smaller focus

issues:
vacuum issues in the MPC (reason: open roughing valve)
FSSS (hall sensor broken - exchange of the entire device planned for tomorrow)

ELOG V3.1.4-7c3fd00